OpenStudy (anonymous):

What is the equation of the line that passes through the points (-1, 3) and (1, 11)? 4x - y = 1 4x + y = 1 x - 4y = -13 4x - y = -7

hero (hero):

Use y = mx + b: 3 = -m + b 11 = m + b Combine the equations together to get 14 = 2b Divide both sides by 2 to get 7 = b Substitute b back into the original equation to get m: 11 = m + 7 11 - 7 = m 4 = m Therefore the equation of the line is y = 4x + 7

OpenStudy (kainui):

Well I'll help you out, basically we have have the formula y=mx+b. m is the slope of the line and b is our y-intercept. So to best figure this out, take your graphing calculator and look at lines of y=x, y=2x, and y=(1/2)x. Play around and discover why slope is described as \[\frac{ rise }{ run }\]

hero (hero):

or you can manipulate that to get 4x - y = -7

OpenStudy (kainui):

What happens when x is 0? Then y=mx+b just sets y=b. So this is why it's called the y-intercept, since at x=0 you're on the y-axis.